OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. The Johnson County Carnival costs $15 plus $0.50 per ride. The Jefferson County Carnival costs $10 plus $1.00 per ride. How many rides should you plan on riding to make the Johnson County Carnival less expensive than the Jefferson County Carnival?

OpenStudy (anonymous):

this is fairly simple to solve jsut by setting up two equations. first, the cost of the johnson carnival is C=15+.5*x where C is the cost and x is the number of rides you go on. now for the jefferson carnival, your equation is C=10+x with the same variables. now we set the two equations equal to one another to see how many rides will make the two equal as any number of rides over this amount will make the johnson carnival less expensive and any number of rides under this amount will make the jefferson carnival less expensive 10+x=15+.5*x now we just solve for x .5*x=5 x=10 so if you take any more than 10 rides at the johnson carnival, you end up spending less than you would at the jefferson carnival